# Quasi-universal relations in the context of future neutron star detections

### v1

@@ III.3.1 Moment of inertia extraction
Figure 5: Moment of inertia distribution from parameter extraction operated with data-driven relations and EoS inference.
We show the moment of inertia distribution extracted with the EoS dependent relation $M(I)$ from the GP+astro and MM+$\chi$+PSR EoSs inferred on by the simulated STROBE-X source in Fig. 4(a). We compare it to the moment of inertia distribution obtained with the data-driven relation fits which parameters are presented in Table 1. The EoS inference leads to a better determination of the moment of inertia than the use of data-driven relation fits. In the case of the GP+astro set, the improvement is significant, while in the case of the meta-model it is almost equivalent: this is in accordance with results presented in Fig 4. The peaks of the EoS inference distributions are not the same for the GP+astro and the meta-model set. The GP+astro set is informed by GW170817 which softens the EoS distribution in that set while the meta-model set can explore stiffer models. The distribution for $I$ determined by the data-driven relation of the GP+astro however explores higher values of the moment of inertia than its EoS inference counterpart. The fit established from a set of EoS informed by data (astrophysics or nuclear physics) does not keep memory distribution of that set: the fit only cares about the boundary points, not the distribution of points itself.
The moment of inertia distribution extracted with gravitational wave data simulated with Cosmic Explorer and Einstein Telescope sensitivity is presented in Fig. 4(b). The extraction of $I$ with the data-driven relation $\bar{I}(\Lambda)$ are very similar: as discussed in Fig. 0(c), this relation is very universal, although we can better distinguish between the two different EoS sets. The difference of the peak’s values in the EoS inference and data-driven relation distributions is also visible even though the distributions overlap well: the use of fits which loose prior information on the sets leads again to higher values of the moment of inertia, suggesting a slightly stiffer EoS. Finally, we also note that the 3G gravitational wave detection offer a better constraint on the moment of inertia than STROBE-X does with EoS inference.
@@ III.3.2 Radius extraction
Figure 6: Radius distribution from parameter extraction operated with data-driven relations and EoS inference.
We show the radius distribution extracted with the EoS dependent relation $C(\Lambda)$ from the GP+astro and MM+$\chi$+PSR EoSs inferred on by the simulated gravitational wave data source in Fig. 6. Results are similar to Fig. 4(b), except that the distribution offered by data-driven relations is much larger. We also note that the meta model inference peak on the radius in Fig 6 is much more pronounced than the peak in moment of inertia of Fig. 4(b). This is because the radius is strongly related to the description of the crust while the moment of inertia is mainly sensitive to the core. While the core treatment of the MM+$\chi$+PSR and the GP+astro seem to lead to similar distributions for the moment of inertia inference, this is not the case for the radius inference. The power of the low density treatment of the meta model explicitly shows in this case.
We also note that the distributions presented in Fig. 5 and Fig. 6 are not always very well resolved. In this paper, we have used pre-computed EoS sets. They will be an issue for future detectors and a generator of EoS set operating with the inference will be necessary to offer well resolved distributions.
@@ IV Conclusion
In this paper, we discuss three macroscopic parameter relations : $C(\Lambda)$, $\bar{I}(C)$ and $\bar{I}(\Lambda)$. We have studied the use of so called quasi-universal relations in moment of inertia and radius extraction from gravitational wave and X-ray signals with current and future detector sensitivity.
We have used three different sets of EoSs to design the relations and their associated fits. We have shown that for compactness related relations discussed in this paper, the EoS variability lets us distinguish between the different sets of EoS while the $\bar{I}(\Lambda)$ presents a more pronounced universality. To accurately express the EoS set dependence of the relations, we refer to data-driven relation instead of quasi-universal relations.
We have discussed the fit error marginalization in parameter extraction and have shown that different approaches do not impact the result for O4 or NICER sensitivity, but do impact the result for Einstein Telescope and Cosmic Explorer and STROBE-X sensitivity.
Finally, we have shown that using data-driven relations in parameter extraction with the future detector sensitivity will lead to much larger uncertainties on the extracted parameter than direct EoS inference. In overall, when the quasi-universality of data-driven relations is overcome by the sensitivity of the detector and its consequent information on the EoS, using data-driven relations is sub-optimal.
In previous usage, data driven relations were broadly equivalent in parameter extraction when compared to direct EoS inference (for example, in GW170817 [21]). We show here that even if future data-driven relations are conditioned on EoS knowledge with precision measurement, direct usage of those EoSs will be more effective in extracting parameters.
@@ Acknowledgements.
The authors thank Rory Smith, Sunny Ng, Philippe Landry, Carl-Johan Haster, Isaac Legred and Emily Wuchner for useful discussion about numerical applications. The authors thank collaborators of the Nuclear Physics from Multi-Messenger Mergers (NP3M) Focused Research Hub for useful discussions. The authors acknowledge the financial support of the National Science Foundation grant Number PHY 21-16686 and PHY 21-10441. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. Software: This work makes use of scipy [52], numpy [53] and matplotlib [54].
Appendix A Mass, radius, moment of inertia and tidal deformability modeling
We study the astrophysical features of a non-rotating neutron star within the gravity theory of General Relativity, assuming the line element ${\rm d}s$ of a spherically symmetric, static and isotropic space-time determined by the metric $g_{\mu\nu}$ as
@@ Appendix A Mass, radius, moment of inertia and tidal deformability modeling
with $r$ the Schwarzschild like radial coordinate. Considering vacuum outside the neutron star, this metric reduces to the Schwarzschild one. Inside the star, the functions $\phi$ (gravitational redshift) and $\lambda$ (radial gravitational distortion) are solved using hydrostatic equilibrium equations and the Equation of State (EoS) ${P=P(\epsilon)}$, with $P$ the pressure and $\epsilon$ the energy density. In the assumption that the matter inside the neutron star is a perfect fluid, we obtain the Tolman-Oppenheimer-Volkov (TOV) differential equations
@@
with $G$ the gravitational constant, $c$ the speed of light and $m(r)$ the gravitational mass radial profile. The central energy density or central pressure, used as a boundary condition to solve the TOV equations, determines the total gravitational mass $M$ and the total radius $R$ of the neutron star.
The dimensionless tidal deformability expresses the quadruple moment response induced on the neutron star by an external gravitational field. Its value at the surface of the star denoted $\Lambda=\lambda_{2}(R)$ is given by the tidal deformability radial profile
@@
with the boundary condition $y(0)=2$, see Section IV.A of Ref. [56]. The functions $F(r)$ and $Q(r)$ are given by
@@
To model the moment of inertia, we use the slow and rigid rotation approximation detailed in Ref. [19] [^6] Modeling the moment of inertia outside of this approximation requires solving the Einstein’s equations with a metric describing a stationary and axisymmetric space-time, see, e.g., the LORENE library.. When considering rigid rotation, the uniform angular frequency is contributed to by the local spin frequency $\omega(r)$ and the angular momentum $j(r)$ of a sphere of radius $r$. We denote $\Omega$ the uniform angular frequency of the star, defined at the surface (${r=R}$) as
@@
The uniform angular frequency, the angular momentum and the moment of inertia are solution of the differential equations
@@
Appendix B Distribution of EoSs in the sets
@@
Appendix C Fit parameters and errors associated to the fits
We present in Table 1 the parameters of the fit for the relations presented in Section II.3.2 and the maximum error associated to the fits with the different EoS sets presented in this paper.
@@ Appendix C Fit parameters and errors associated to the fits
Table 1: Fit parameters of the relation $C(\Lambda)$, $\bar{I}(C)$ and $\bar{I}(\Lambda)$ based on the parametrization presented in Eqs. (1), (2), (3), and maximum error associated to the fit $\Delta_{\rm max}$. Results are presented for the EoS sets discussed in this paper, GP+astro, the nuclear set and the MM+$\chi$+PSR.

### v2

@@ III.3.1 Moment of inertia extraction
Figure 5: Moment of inertia distribution from parameter extraction operated with quasiuniversal relation fits and EoS inference.
We show the moment of inertia distribution extracted from the simulated STROBE-X source in Fig. 5(a). We perform EoS inference with either the GP+astro and MM+$\chi$+PSR EoS sets as priors, and compute the resulting $I$ distribution using the updated EoSs and the observed masses. We compare this to the moment of inertia distribution obtained with the quasiuniversal relation fits which parameters are presented in Table 1. The EoS inference leads to a better determination of the moment of inertia than the use of quasiuniversal relation fits. In the case of the GP+astro set, the improvement is significant, while in the case of the metamodel it is almost equivalent: this is in accordance with results presented in Fig 4. The peaks of the EoS inference distributions depend on the EoS prior and are not the same for the GP+astro and the metamodel. The GP+astro set is informed by GW170817, which softens the EoS distribution, while the metamodel set allows stiffer models. The distribution for $I$ determined by the quasiuniversal relation fits with the GP+astro extracts higher values of the moment of inertia than its EoS inference counterpart. This is because the quasiuniversal fits discard some of the prior information encoded in the full EoS set; it considers only the relation between compactness and $I$ rather than the preferred values of compactness within that relation, extracting values linked only to the current observation.
The moment of inertia distribution extracted with gravitational wave data simulated with CE and ET sensitivity is presented in Fig. 5(b). The extraction of $I$ with the quasiuniversal relation fits of $\bar{I}(\Lambda)$ are very similar: as discussed in Fig. 1(c), this relation is very universal. The difference of the peak’s values in the distributions for EoS inference and quasiuniversal relation fits is also visible even though the distributions overlap well; both EoS sets generate similar fit results, as those discard prior information preferring softer or stiffer EoS. Finally, we also note that the 3G gravitational wave detection offer a better constraint on the moment of inertia than STROBE-X does with EoS inference.
@@ III.3.2 Radius extraction
Figure 6: Radius distribution from parameter extraction for simulated CE+ET gravitational wave data, using quasiuniversal relation fits and EoS inference.
We show the radius distribution extracted with the fit of the quasiuniversal relation fit of $C(\Lambda)$ from the GP+astro and MM+$\chi$+PSR EoSs, and from the updated EoS distributions inferred from the simulated gravitational wave data source in Fig. 6. Results are similar to Fig. 5(b), except that the distribution offered by quasiuniversal relation fit is much larger. The extraction of the radius from the GP+astro set leads to a range extending to very large radii. The systematic bias due to the use of fit in the extraction of the radius was also discussed in Ref. [61, 62]. We note that the metamodel inference peak on the radius in Fig 6 is much narrower than the peak in moment of inertia of Fig. 5(b). This is because the radius is strongly related to the description of the crust while the moment of inertia is mainly sensitive to the core. While the core treatment of the MM+$\chi$+PSR and the GP+astro seems to lead to similar distributions for the moment of inertia inference, this is not the case for the radius inference. The power of the low density treatment of the meta model explicitly shows in this case. This shows the impact of the nuclear physics information in the MM+$\chi$+PSR, which limits the EoS range as seen in Fig. 7.
We note that the distributions presented in Fig. 5 and Fig. 6 are not always very well resolved. In this paper, we have used precomputed EoS sets as priors for inference following methods described in Refs. [28, 52, 63]. This will be an issue for future detectors; a generative model for additional draws from the EoS prior operating with the inference would be necessary to offer well-resolved distributions.
We also note that for this work we have considered a single measurement that offers one constraint on the extracted parameter. If multiple measured parameters allow extraction with more than one technique, then it would be possible to find tension in the extracted values, as discussed in Ref. [64]. Such tension indicates that the prior assumptions of the EoS set have been violated, or in the case of a quasiuniversal extraction that the true EoS has a relation outside the fit error, and could suggest currently-unmodeled features in the EoS.
@@ IV Conclusion
In this paper, we discuss three macroscopic quasiuniversal relations : $C(\Lambda)$, $\bar{I}(C)$, and $\bar{I}(\Lambda)$. We have studied the use of these quasiuniversal relations in moment of inertia and radius extraction from gravitational wave and x-ray signals with current and future detector sensitivity.
We have used three different sets of EoSs to calibrate the relations and their associated error distributions. We have shown that for $C(\Lambda)$ and $\bar{I}(C)$ relations, the variability of EoS in these sets leads to differences in their corresponding quasiuniversal fit, quantified here by a larger fit error, while the $\bar{I}(\Lambda)$ presents a more pronounced universality.
We have discussed the fit error marginalization in parameter extraction and have shown that different approaches do not significantly impact the result for O4 or NICER sensitivity, but do impact the result for Einstein Telescope and Cosmic Explorer and STROBE-X sensitivity.
Finally, we have shown that using quasiuniversal relation fits in parameter extraction with the future detector sensitivity will overestimate uncertainties on the extracted parameter when compared to direct EoS inference. In general, when the quasiuniversality fit error gives larger variation in the extracted parameter than the uncertainty on that parameter from the measurement error, using quasiuniversal relation fits no longer reflects the information gained about that parameter from the observation and its EoS implications.
At current observational precision, quasiuniversal relations give broadly equivalent results for parameter extraction when compared to direct EoS inference for the same quantities. We show here that quasiuniversal relations become less effective when projecting results for future observations, even when conditioned on EoS knowledge from current precision measurement. Instead, the direct EoS constraints inferred from future observations should be included when determining observational implications to reveal the full capability of next-generation facilities.
@@ Acknowledgements.
The authors thank Rory Smith, Sunny Ng, Philippe Landry, Carl-Johan Haster, Isaac Legred and Emily Wuchner for useful discussion about numerical applications. The authors thank collaborators of the Nuclear Physics from multimessenger Mergers (NP3M) Focused Research Hub for useful discussions. The authors acknowledge the financial support of the National Science Foundation Grants No. PHY 21-16686 and No. PHY 21-10441. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants No. PHY-0757058 and No. PHY-0823459. Software: This work makes use of scipy [65], numpy [66] and matplotlib [67].
Appendix A Equation of state relation for agnostic sets
Figure 7 presents the equation of state (relation between the pressure $P$ and the baryonic density $n_{B}$) for the two agnostic sets discussed in this paper.
Figure 7: Contours (99 percentile) for the relation between the pressure and the baryonic density for the GP+astro and MM+$\chi$+PSR agnostic sets.
Appendix B Mass, radius, moment of inertia and tidal deformability modeling
We study the astrophysical features of a nonrotating neutron star within the gravity theory of general relativity, assuming the line element ${\rm d}s$ of a spherically symmetric, static and isotropic space-time determined by the metric $g_{\mu\nu}$ as
@@ Appendix B Mass, radius, moment of inertia and tidal deformability modeling
with $r$ the Schwarzschild-like radial coordinate. Considering a vacuum outside the neutron star, this metric reduces to the Schwarzschild one. Inside the star, the functions $\phi$ (gravitational redshift) and $\lambda$ (radial gravitational distortion) are solved using hydrostatic equilibrium equations and the EoS ${P=P(\epsilon)}$, with $P$ the pressure and $\epsilon$ the energy density. In the assumption that the matter inside the neutron star is a perfect fluid, we obtain the TOV differential equations
@@
with $G$ the gravitational constant, $c$ the speed of light, and $m(r)$ the gravitational mass radial profile. The central energy density or central pressure, used as a boundary condition to solve the TOV equations, determines the total gravitational mass $M$ and the total radius $R$ of the neutron star.
The dimensionless tidal deformability expresses the quadruple moment response induced on the neutron star by an external gravitational field. Its value at the surface of the star, denoted $\Lambda=\lambda_{2}(R)$, is given by the tidal deformability radial profile,
@@
with the boundary condition $y(0)=2$, see Sec. IV.A of Ref. [69]. The functions $F(r)$ and $Q(r)$ are given by
@@
To model the moment of inertia, we use the slow and rigid rotation approximation detailed in Ref. [31] [^5] Modeling the moment of inertia outside of this approximation requires solving the Einstein’s equations with a metric describing a stationary and axisymmetric space-time, see, e.g., the LORENE library.. When considering rigid rotation, the uniform angular frequency is contributed to by the local spin frequency $\omega(r)$ and the angular momentum $j(r)$ of a sphere of radius $r$. We denote $\Omega$, the uniform angular frequency of the star, defined at the surface (${r=R}$) as
@@
The uniform angular frequency, the angular momentum and the moment of inertia are solutions of the differential equations,
@@
Appendix C Distribution of EoSs in the sets
@@
Appendix D Fit parameters and errors associated to the fits
We present in Table 1 the parameters of the fit for the relations presented in Sec. II.3.2 and the maximum error associated to the fits with the different EoS sets presented in this paper.
@@ Appendix D Fit parameters and errors associated to the fits
Table 1: Fit parameters of the relation $C(\Lambda)$, $\bar{I}(C)$, and $\bar{I}(\Lambda)$ based on the parametrization presented in Eqs. (1), (2), (3), and maximum error associated to the fit $\Delta_{\rm max}$. Results are presented for the EoS sets discussed in this paper, GP+astro, the nuclear set and the MM+$\chi$+PSR.